OpenStudy (anonymous):

What is the value of y when x = -1? 5x + 4y = 11 -4 -3 3 4

OpenStudy (anonymous):

5x + 4y = 11 5(-1) + 4y = 11 -5 + 4y = 11 4y = 16 y = 4